a photon of 300nm is absorbed by a gas and then re emitts two photons.one reemitted photon has a wavelength of 500nm.calculate energy of other photon

Grade:11

1 Answers

Umakant biswal
7 years ago
@ roohinaya 
i am giving you the correct approach to solve this question . 
energy absorbed for photon = hc / lambda (1 ) = hc / 300 * 10 ^ -9 
energy reemitted out for phtoton = hc / lambda (2) = hc / 500 * 10 ^-9 
energy absorbed = energy for emitted photon = energy for 2nd emitted photon 
energy for 2nd emitted photon = hc / 500 * 10 ^ -9 
now put the value of h as planks constant and c as the speed of light , 
then u will get the answer . 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..